

********************************************************************
UARS WEEKLY STATUS REPORT - 6 JULY 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM was turned off for most of this week. Turn on is scheduled for
Friday. Until turnoff SUSIM successfully gathered its normal solar,
offset, and occultation scan data without apparent problems.
SUSIM Version 7 production software is in Engineering Mode testing
at the CDHF.
***********
PEM
***********
PEM/AXIS was active from June 22 through June 27. During this time,
PEM/AXIS generated energy deposition from inverted X-ray spectra for each of
its 16 look directions.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS has had a recurrence of antenna scan slips, similar to what it experienced
several years ago and eliminated then by changing the method of instrument
operation. The recent slips were detected, and corrected, by on-board software
before they caused any damage to the instrument or had any impact on its
science data. MLS is now being operated in a manner to obtain diagnostic
information on the scan slips, or else is put in safehold while the engineering
data are being analyzed. Initial analyses indicate no change in play in the
scan mechanism, and that the scan motor performance is unchanged since the
start of mission. Additional tests are being conducted to, hopefully, determine
an acceptable operating mode.
***********
HRDI
***********
HRDI is currently off. Next scheduled turn-on is July 3rd.
Mark Burrage will be presenting a review paper, "Mesosphere and lower
thermosphere coupling from UARS/HRDI," at the COSPAR meeting in Nagoya,
Japan, July 12-19.
***********
WINDII
***********
WINDII is preparing for another PMC observing session later this month.
Final touches are also being made to Version 5 of the production
processing software.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
An orbit-adjust maneuver for the UARS spacecraft is being planned
for 12 August 1998. In preparation for the thruster firing on
12 August, small thruster "burns" for calibration are being
planned for 28 July 1998.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
28 days HAL (4.80) PROD (L1) (1868R-539C-8F=1321L)
0 days HAL (4.80) PROD (L3) (1868R-306C=1562L)
347 days SOL (4.61) PROD (L1) (2922R-2922C=0L)
120 days SOL (4.61) PROD (L2-L3) (2349R-687C=1662L)
The CDHF received SUSIM V7.0 Engineering Software Delivery.
The CDHF completed the 9-track tape to 8mm tape conversion for
the Level 0 Catalog Saves stored at TSSF. UARS days 4 through 1175 were
converted. 586 8mm tapes replaced 3592 9-track tapes.
The CDHF will load May 1998 current day DAAC data in July 1998.
CDHF pending software upgrades: IDL 5.1, OpenIngres, Fortran V6.5.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 June 1998 (Orbit 36903, GMT Day 164) through 19 June
1998 (Orbit 37006, GMT Day 170).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 171/0047 37008 On
WINDII 172/2313 37037 On
WINDII 173/0056 37038 Off
WINDII 173/1156 37044 On
PEM 174/1019 37058 AXIS on
The beta angle reached a minimum angle of 0.0 degrees on 15 June 1998 (DOY
166). The beta angle is now increasing toward a maximum angle of 33.8
degrees on 30 June 1998 (DOY 181).
A series of SSPP gridded offset maneuvers requested by SUSIM operations was
performed on GMT Days 171 through 176. The offsets utilized the 5 and 10
arcminute intervals (as opposed to 15 and 30 arcminute intervals) with two
identical offsets being performed on consecutive orbits of each day.
Another series of similar gridded offsets is planned for GMT Days 185
through 190.
A Rmanually generatedS stellar offset experiment requested by SOLSTICE
operations was performed on GMT day 176.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.5 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.1 degrees C. Additional battery performance data for
each battery on 20 June 1998 and 26 June 1998 are provided in Appendix D.
The clock error ranged from -10.7 to 10.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 10
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HALOE 03
% HRDI 06
% PEM 01
% WINDII 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 05
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/176 show the observatory data loss to be
20 hours, 44 minutes, 38.388 seconds (an increase of 32.488 seconds since
last report period). This is a 0.0350 percent data loss which equals a
99.9650 percent data capture for the mission.
The increase of 32.488 seconds occurred over a two day period:
7 23.104 seconds on Day 171 due to insufficient overlap, and
7 9.344 seconds on Day 173 due to insufficient overlap.
The latest version of the re-engineering teamUs quicklook generation
software was tested on UMPGAL. Problems still exist in the flagging of bad
quality frames.
Version 2.33 of the UGPL software for CDHF/UMPGAL was released on 22 June
1998. This version has larger telemetry point index tables (increased from
6000 points to 7500 points). This change was required to support the new
telemetry generated by the new flight software (V63).
Testing the command side of the I/P Net remains on hold.
OTHER SIGNIFICANT EVENTS
Continued the upgrade to the OBC Power Monitor (PMON) flight software.
Ground System testing continued this week. Several builds of the ground
system including 10.07 level, 10.08 level, and development packs
demonstrated the same failure associated with the dump directive. So far
only Pack PE53 has shown the desired behavior and operated properly in the
absence of default parameters.
Work continues on the UARS orbit adjust planning re-scheduled for 12 August
1998. Additionally, planning is underway for the 28 July 1998 thruster
tests (engineering test).
Representative plots of UARS historical power subsystem data were
accumulated and provided to EOS-AM spacecraft personnel. The plots contain
information pertaining to the modular power subsystem, main and auxiliary
solar arrays and beta angle for beta values between 10 and 20 degrees.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
A meeting was held on 26 June between FOT members and CSC personnel (Sidney
Kirschner) to discuss Year 2000 related testing of the UARS CMS. Mr.
Kirschner will be working closely with the FOT over the next several months
to ensure the CMS is thoroughly tested for Year 2000 compatibility.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 15 July 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 12 August 1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 June 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37008 171/0047 98-204 HALOE Power ON
37012 171/0747 98-205 HRDI Heater Cycle to 150 Amp/Mins
37013 171/0925 98-206 HRDI Heater Cycle to 175 Amp/Mins
37014 171/1149 98-207 HRDI Low PWR Mode
37019 171/1843 98-208 HALOE Power ON/Disable MUB
37020 171/2028 98-209 HALOE Power OFF/ON
37031 172/1452 98-210 HRDI Heater Delay from 175 to 150 A/M
37037 172/2313 98-211 HRDI Heater Delay from 150 to 125 A/M
37037 172/2313 98-212 WINDII Power ON
37038 173/0056 98-213 WINDII Power OFF
37044 173/1156 98-214 WINDII Power ON
37055 174/0519 98-215 HRDI Heater Delay to 100 amp/mins
37058 174/1019 98-216 PEM Power ON AXIS
37059 174/1200 98-217 OBC Clk Adj fr 0.00 to 20.6 msec/day
37079 175/2003 98-218 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 June 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HAH28VA R-LO Heater Motor Voltage 37011-19
HRIFEELECTMP R-HI IFE Electronics Temp 37013-14
HRIFESURVTMP Y-HI IFE Survival Heater Temp 37013-14
HRRADTEMP Y-HI IFE Radiator Temp 37013-14
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 37026-33,42-48,55-
64,70-72,86,117,121
WIEUTEMP Y-LO WINDII EU Temperature 37008-44
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 37015-112
HRIFESURVTMP R-HI IFE Survival Heater Temp 37013-14
HASMIRRT Y-HI HALOE Sec. Mirror Temp 37018-19
WIQBCURRENT Y-LO WINDII Quiet Bus. Current 37044
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 June 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37059 174/12:14 (23 Jun 98) from 00.0 to 20.6 msec/day
37079 175/20:12 (24 Jun 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 June 1998
20 Jun 1998 - GMT Day 171 Beta = 14.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.2 0.000 0.0
2 +22.4/ -22.4 +7.09/+4.51 +31.2/-16.4 23.5 1.020 23.3
3 +72.8/ -5.6 +5.47/+3.23 +30.4/-17.6 23.5 1.021 22.6
26 Jun 1998 - GMT Day 177 Beta = 29.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.2 0.000 0.0
2 +22.4/ -22.4 +7.09/+5.47 +30.4/-16.0 24.5 1.022 23.3
3 +39.2/ -5.6 +5.79/+4.51 +29.6/-16.4 24.5 1.023 22.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 June 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-028 FOT 37039 173/0259 14170 Generic Late Acquisition
98-033 FOT 37080 175/2143 N/A TAC #2 Crashed
98-035 FOT 37083 176/0129 N/A TAC #2 Crashed
98-036 FOT 37083 176/0844 N/A TAC #2 Crashed
98-038 FOT 37098 177/0134 N/A TAC #8 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-030 FOT 36910 164/12:50 TAC 6 CRASH
98-031 FOT 36930 165/19:55 TAC 6 CRASH
98-032 FOT 36973 168/16:51 TAC 4 CRASH
98-033 FOT 37080 175/21:55 TAC 2 CRASH
98-034 FOT 37010 171/04:22 TMON 1 TRIP
98-035 FOT 37083 176/01:29 TAC 2 CRASH
98-036 FOT 37087 176/08:44 TAC 2 CRASH
98-037 FOT 36887 163/00:02 19991 XMTR OFF AT AOS
98-038 FOT 377098 177/01:34 TAC 2 CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-028 FOT SEE LIST JUNE 1998 14170 GENERIC LATE ACQ